□ Emily has decided to spend one-third of her income on clothing. What is Emily’s demand curve for clothing?




□ Emily’s taste – as reported in the previous question - has now changed. She now realizes the importance of savings and spends only one-tenth of her income on clothing, instead of one-third. How will Emily’s demand curve change?


1
Expert's answer
2021-07-28T06:37:02-0400

a)demand function

"Q = a - bP"

where Q=linear demand curve,a=factors influencing demand besides price,b=slope and p=price

"Q = a - \frac{1}{3}P"

b)demand function

"Q = a - bP"

where Q=linear demand curve,a=factors influencing demand besides price,b=slope and p=price

"Q = a - \frac{1}{10}P"